Child pages
  • Commander Report Day 07 TasMars2013
Skip to end of metadata
Go to start of metadata

Date Time: 20:00 02/01/2013
Written By: David Willson, Crew Commander of MDRS123

Today was a quiet day spent mostly reviewing the results of the previous day’s EVA.

The temperature today did not exceed 50° F, the sky cloud free.

Jen and David reviewed and analyzed the Raman spectra taken from the previous day field trip. They undertook a short 1 hour EVA at 4.30 pm to location USNG 15SUD0370514711 collecting cretaceous sand and silt mud samples. David also prepared the quadcopter assembly for flight.

Melanie trialed two recipes and completed and sent an interim report on Mars exploration recipes, using long term storable food. Melanie also reviewed with the crew the planetary protection risk analysis work that will be undertaken during next week.

Emma installed a fan into the roof vent and we noticed a clear improvement in air quality. This was a qualitative assessment. Emma also itemized the medicine cabinet.

The arrival of the port-a-loo did not occur as expected by Mission Support.


  • No labels